# Soft Reasoning Paths for Knowledge Graph Completion

### v1

@@ Abstract
Reasoning paths are reliable information in knowledge graph completion (KGC) in which algorithms can find strong clues of the actual relation between entities. However, in real-world applications, it is difficult to guarantee that computationally affordable paths exist toward all candidate entities. According to our observation, the prediction accuracy drops significantly when paths are absent. To make the proposed algorithm more stable against the missing path circumstances, we introduce soft reasoning paths. Concretely, a specific learnable latent path embedding is concatenated to each relation to help better model the characteristics of the corresponding paths. The combination of the relation and the corresponding learnable embedding is termed a soft path in our paper. By aligning the soft paths with the reasoning paths, a learnable embedding is guided to learn a generalized path representation of the corresponding relation. In addition, we introduce a hierarchical ranking strategy to make full use of information about the entity, relation, path, and soft path to help improve both the efficiency and accuracy of the model. Extensive experimental results illustrate that our algorithm outperforms the compared state-of-the-art algorithms by a notable margin. The code will be made publicly available after the paper is officially accepted.
@@ 2.1 Knowledge Graph Completion
Existing methods can be broadly classified into two categories: embedding-based methods and text-based methods. Embedding-based methods represent entities and relations as vectors. Translational models, such as TransE Bordes et al. (2013) and TransH Wang et al. (2014), are computationally efficient but struggle to capture complex relationships. Tensor decomposition models like ComplEx Trouillon et al. (2016) are effective at modeling diverse patterns but face scalability challenges. Graph neural networks, such as CompGCN Vashishth et al. (2019), enhance entity and relation representations by incorporating neighborhood information, though they require careful design to handle complex data structures. Text-based methods utilize textual descriptions or context associated with entities and relations to perform knowledge graph completion (KGC). For instance, KG-BERT Yao et al. (2021) employs BERT to encode triples as text sequences and classifies their validity. SimKGC Wang et al. (2022a) and C-LMKE Wang et al. (2022b) propose a contrastive learning framework to build more discriminative knowledge graph models. Despite their potential, text-based methods typically determine the target triple based solely on the similarity between the pair $(h,r)$ and candidate entity text embeddings, often overlooking other valuable auxiliary information.
@@ 2.2 Reasoning path in KGC
In knowledge graph completion, reasoning plays a pivotal role in accurately predicting missing links or facts. While traditional approaches typically rely on direct embeddings of entities and relations, reasoning path-based methods enhance the model’s capacity to capture higher-order relationships between entities. These methods focus on identifying paths through the graph that may represent logical or semantic relations, thereby guiding the completion process. GraIL Teru et al. (2020) utilizes GNNs to evaluate the relevance between reasoning path relations and the target relation, while BERTRL Zha et al. (2021) employs BERT to encode reasoning paths and candidate triples to compute their scores. ReDistLP Iwamoto and Kameiwa (2024) completes the task by aggregating information from various reasoning paths. These methods have demonstrated strong performance in inductive KGC tasks. However, they face challenges when dealing with situations involving missing reasoning paths and a large number of candidate triples.
@@ 2.3 Prompt tuning
Prompt tuning, through the use of prompts, enables pre-trained language models (PLMs) to achieve exceptional performance across various downstream tasks with minimal computational cost. CSProm-KG Chen et al. (2023) is the first work to incorporate prompt tuning into KGC tasks. By applying prefix tuning in conjunction with GNNs, it effectively completes the KGC task under low-parameter conditions. AutoKG Zhu et al. (2023) also explores the application of prompt engineering within the knowledge graph domain. A frequently overlooked aspect of prompt tuning is its capacity to learn general representations of data during the training process, a feature that our method leverages. This enables our model not only to handle specific tasks but also to extract and utilize general patterns from the data, thereby enhancing the model’s generalization ability and overall performance.
@@ 4.2 Performance Comparison with SOTA Method
In this study, we conducted a comparative analysis of SRP-KGC, comparing it with both embedding-based and text-based approaches. The embedding-based methods include TransE Bordes et al. (2013), ComplEx Trouillon et al. (2016), RotatE Sun et al. (2019b), ConvE Dettmers et al. (2017), CompGCN Vashishth et al. (2019), TuckER Balazevic et al. (2019), CompoundE Ge et al. (2023), KRACL Tan et al. (2022), and RotatE-VLP Li et al. (2023). On the other hand, the text-based methods include KG-BERT Yao et al. (2021), StAR Wang et al. (2021), KG-S2S Chen et al. (2022), C-LMKE Wang et al. (2022b), SimKGC Wang et al. (2022a), CSProm-KG Chen et al. (2023), LP-BERT Li et al. (2022), GS-KGC Yang et al. (2024) and GHN Qiao et al. (2023).
@@
Table 4: The Hits@10 of FB15k-237 under inductive scenario. The optimal values of each metric are marked in blod.
@@ A.2 Inductive datasets
A.3 The effectiveness of hierarchical ranking.

### v2

@@ Abstract
Reasoning paths are reliable information in knowledge graph completion (KGC) in which algorithms can find strong clues of the actual relation between entities. However, in real-world applications, it is difficult to guarantee that computationally affordable paths exist toward all candidate entities. According to our observation, the prediction accuracy drops significantly when paths are absent. To make the proposed algorithm more stable against the missing path circumstances, we introduce soft reasoning paths. Concretely, a specific learnable latent path embedding is concatenated to each relation to help better model the characteristics of the corresponding paths. The combination of the relation and the corresponding learnable embedding is termed a soft path in our paper. By aligning the soft paths with the reasoning paths, a learnable embedding is guided to learn a generalized path representation of the corresponding relation. In addition, we introduce a hierarchical ranking strategy to make full use of information about the entity, relation, path, and soft path to help improve both the efficiency and accuracy of the model. Extensive experimental results illustrate that our algorithm outperforms the compared state-of-the-art algorithms by a notable margin. Our code will be released at https://github.com/7HHHHH/SRP-KGC.
@@ 2.1 Knowledge Graph Completion
Existing methods for knowledge graph completion (KGC) fall into two categories: embedding-based and text-based.Embedding-based methods encode entities and relations as vectors. Translational models (e.g., TransE Bordes et al. (2013), TransH Wang et al. (2014)) are efficient but weak in modeling complex patterns. Tensor models like ComplEx Trouillon et al. (2016) handle diverse relations but scale poorly. Graph neural networks (e.g., CompGCN Vashishth et al. (2019)) incorporate neighbor information to improve representations, though they require careful architecture design.Text-based methods leverage textual context. KG-BERT Yao et al. (2021) encodes triples as text for classification. SimKGC Wang et al. (2022a) and C-LMKE Wang et al. (2022b) use contrastive learning for better discrimination. However, these methods often rely only on $(h,r)$ and candidate entity text similarity, ignoring richer auxiliary cues.
@@ 2.2 Reasoning Path in KGC
Reasoning is crucial for accurate knowledge graph completion (KGC). Unlike traditional embedding-based methods, reasoning path-based approaches capture higher-order relations by exploring paths that reflect semantic or logical connections. GraIL Teru et al. (2020) uses GNNs to assess path-relation relevance, BERTRL Zha et al. (2021) encodes reasoning paths and candidate triples with BERT, and ReDistLP Iwamoto and Kameiwa (2024) aggregates multiple paths for prediction. These methods excel in inductive KGC but struggle when reasoning paths are missing or candidate triples are abundant.
@@ 2.3 Prompt Tuning
Prompt tuning Hou et al. (2024), through the use of prompts, enables pre-trained language models (PLMs) Hou et al. (2025) to achieve exceptional performance across various downstream tasks with minimal computational cost. CSProm-KG Chen et al. (2023) is the first work to incorporate prompt tuning into KGC tasks. By applying prefix tuning in conjunction with GNNs, it effectively completes the KGC task under low-parameter conditions. AutoKG Zhu et al. (2023) also explores the application of prompt engineering within the knowledge graph domain. A frequently overlooked aspect of prompt tuning is its capacity to learn general representations of data during the training process, a feature that our method leverages. This enables our model not only to handle specific tasks but also to extract and utilize general patterns from the data, thereby enhancing the model’s generalization ability and overall performance.
@@ 4.2 Performance Comparison with SOTA Method
In this study, we conducted a comparative analysis of SRP-KGC, comparing it with both embedding-based and text-based approaches. The embedding-based methods include TransE Bordes et al. (2013), ComplEx Trouillon et al. (2016), RotatE Sun et al. (2019b), ConvE Dettmers et al. (2017), TuckER Balazevic et al. (2019), CompoundE Ge et al. (2023), KRACL Tan et al. (2022), and RotatE-VLP Li et al. (2023). On the other hand, the text-based methods include KG-BERT Yao et al. (2021), StAR Wang et al. (2021), KG-S2S Chen et al. (2022), C-LMKE Wang et al. (2022b), SimKGC Wang et al. (2022a), CSProm-KG Chen et al. (2023), LP-BERT Li et al. (2022), GS-KGC Yang et al. (2024) and GHN Qiao et al. (2023).
@@
Table 4: The Hits@10 of FB15k-237 under the inductive scenario. The optimal values of each metric are marked in bold.
@@ 6 Conclusion
Acknowledgments
This work was supported by the Huxiang Young Talents in Science and Technology Innovation Project (No. 2024RC3148).
@@ A.2 Inductive datasets
A.3 The effectiveness of hierarchical ranking